OpenStudy (anonymous):

classify the solutions of 1/x+4+1/2=1/x+4

OpenStudy (anonymous):

@Ashleyisakitty @hummel650

OpenStudy (anonymous):

@twvogels

OpenStudy (anonymous):

@zubhanwc3

OpenStudy (anonymous):

@samarasamara

OpenStudy (anonymous):

@waheguru

OpenStudy (anonymous):

@mahone331

OpenStudy (anonymous):

are those division signs or fractions?

OpenStudy (anonymous):

fractions!

OpenStudy (anonymous):

@wolfe8 help plz!

OpenStudy (wolfe8):

What does it mean classify the solutions?

OpenStudy (anonymous):

well, shes offline :/

OpenStudy (anonymous):

Im so soorry I was rushing home! @wolfe8

OpenStudy (wolfe8):

It's ok. What does it mean by classify the solutions?

OpenStudy (anonymous):

I guess to simplify..one of the answers is x=-4 extraneous

OpenStudy (anonymous):

@wolfe8

OpenStudy (wolfe8):

I mean, I can solve for x but not sure about classifying...

OpenStudy (anonymous):

Well lets see what it comes out too!

OpenStudy (wolfe8):

Oooh wait you can't solve it normally. It wouldn't have a solution. I can't remember how to do this. I'm sorry. I thought it was a simple one.

OpenStudy (anonymous):

Ok thankss

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

@ganeshie8 please help..

ganeshie8 (ganeshie8):

yes u may use below word to classify : "No solutions"

OpenStudy (anonymous):

its not an option :/

ganeshie8 (ganeshie8):

wat are your options ? :) you may use 'inconsistent' but that word is reserved for a system of equations

OpenStudy (anonymous):

x=-4 extraneous x=-4 non extraneous x=-8 extraneous x=-8 non extraneous

ganeshie8 (ganeshie8):

okay, then we're interpreting the equation incorrectly.

ganeshie8 (ganeshie8):

can u take the snapshot of original q and post it ?

ganeshie8 (ganeshie8):

1/x+4+1/2=1/x+4 can be interpreted in many ways, cuz parenthesis are missing here

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

thank you :) lets solve this.

ganeshie8 (ganeshie8):

multiply both sides wid 2(x+4), you wud get : 2 + x+4 = 2

ganeshie8 (ganeshie8):

simplifying, gives x = -4 however, u cannot put this solution in the equation cuz that wud result in a 0 in the denominator of fraction. so the answer wud be x = -4 and extraneous

ganeshie8 (ganeshie8):

extraneous solution = illegal/ fake solution

OpenStudy (anonymous):

Ugh thank you soo much! !

OpenStudy (anonymous):

Can you helo with one more!

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

get rid of fractions

ganeshie8 (ganeshie8):

multiply 10x both sides

ganeshie8 (ganeshie8):

wat do u get ? :)

OpenStudy (anonymous):

20 and 30?

ganeshie8 (ganeshie8):

is latex working for u ? do u see below in red ? \(\color{red}{\text{do u see this in red ?}}\)

OpenStudy (anonymous):

Wait whaaa?

ganeshie8 (ganeshie8):

for many users latex is not working today. let me knw if u see my previous reply in red color :o

OpenStudy (anonymous):

Im not working for anyone lol

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

latex = equation editer ok

ganeshie8 (ganeshie8):

today 'equation editor' is not working lol

OpenStudy (anonymous):

ohhh hhahaha

ganeshie8 (ganeshie8):

after multiplying 10x both sides u wud get :- 4x + 6 = x(x+5) x^2 + x - 6 = 0 you can solve this now ?

OpenStudy (anonymous):

add 6 on both sides

ganeshie8 (ganeshie8):

nope, we need to factorize it

OpenStudy (anonymous):

Aw man ok

ganeshie8 (ganeshie8):

x^2 + x - 6 = 0 (x+3)(x-2) = 0 x = -3 or x = 2

OpenStudy (anonymous):

Mhmm I see thanks so much . youre the best

ganeshie8 (ganeshie8):

nope, you're the best :) u wlc !